Institute for Language Technology & Artificial Intelligence (ITK)

Tilburg University

Tilburg

the Netherlands

Researchers Peter Flach.

ILP Areas ILP theory, Induction of integrity constraints, Predicate invention.

Description of Research The work on ILP done at ITK concentrates on two main topics: a logical theory of inductive reasoning, and the discovery of regularities in data (weak induction). The first topic aims at providing a unifying framework for different kinds of induction, and at studying the relationships between them, and with other kinds of reasoning. The second topic is directed towards methods for inducing non-classification rules, such as integrity constraints. These two topics are closely related, since a unifying framework for weak induction and the classification-oriented notion of strong induction does not yet exist.

The logical theory of induction under development revolves around the notion of an inductive consequence relation. In [2] a number of useful properties of such a consequence relation is identified. In [5] a minimal set of properties of inductive consequence relations is identified. Consequently, strong and weak induction can be seen as special cases of this canonic type of induction. Current work aims at developing a model-theoretic account of inductive consequence relations.

The work on discovery of regularities in data started with a study of induction of attribute dependencies in relational databases [1]. In cooperation with the Jožef Stefan Institute the algorithm for inducing functional dependencies has been improved considerably [3]. Future work includes improving the multivalued dependency algorithm along the same lines. Building on this work, a prototype system called INDEX has been implemented, which restructures a knowledge base on the basis of induced attribute dependencies. Since the restructured knowledge base will contain predicates not present in the original knowledge base, this also resulted in a new approach to predicate invention through integrity constraints [4].

Recent Publications

[1] P. Flach. Inductive characterisation of database relations. In Z.W. Ras, M. Zemankova & M.L. Emrich (eds.) Proc. Methodologies for Intelligent Systems ISMIS’90, 371–378, North-Holland, Amsterdam, 1990. Full version appeared as ITK Research Report No. 23.

[2] P. Flach. A framework for Inductive Logic Programming. In S. Muggleton (ed.) Inductive Logic Programming, 193–211. Academic Press, London, 1992.

[3] I. Savnik, P. Flach. Bottom-up induction of functional dependencies from relations. In G. Piatetsky-Shapiro (ed.) Proc. AAAI’93 Workshop on Knowledge Discovery in Databases, 174–185, 1993.

[4] P. Flach. Predicate invention in inductive data engineering. In P. Brazdil (ed.) Proc.

European Conference on Machine Learning, ECML’93, Lecture Notes in Artificial Intelligence, Vol. 667, 83–94. Springer-Verlag, 1993.

[5] P. Flach. A model of inductive reasoning. In M. Masuch & L. Polos (eds.), Knowledge Representation and Reasoning under Uncertainty, Springer-Verlag, Berlin, 1994.

Contact Person Peter A. Flach Institute for Language Technology & Artificial Intelligence Tilburg University POBox 90153 5000 LE Tilburg the Netherlands Phone: +31 13 663119 Fax: +31 13 663069 E-mail: flach@kub.nl